OpenStudy (anonymous):

find the sum: 5 6/7 + 3 2/3

OpenStudy (anonymous):

@Jennifer_Thomas

OpenStudy (anonymous):

ok if you want to find the sum you first need to have the same denomianator

OpenStudy (anonymous):

closest is 21

OpenStudy (anonymous):

\[5\frac{ 6 }{ 7 } + 3\frac{ 2 }{ 3 } = 5\frac{ 18 }{ 21 } + 3\frac{ 14 }{ 21 }\] what would you get for your answer

OpenStudy (anonymous):



OpenStudy (anonymous):

that took me a while :P